# Generative Artificial Intelligence for Beamforming in Low-Altitude Economy

### v1

@@ Abstract
The growth of low-altitude economy (LAE) has driven a rising demand for efficient and secure communication. However, conventional beamforming optimization techniques struggle in the complex LAE environments. In this context, generative artificial intelligence (GenAI) methods provide a promising solution. In this article, we first introduce the core concepts of LAE and the roles of beamforming in advanced communication technologies for LAE. We then examine their interrelation, followed by an analysis of the limitations of conventional beamforming methods. Next, we provide an overview of how GenAI methods enhance the process of beamforming, with a focus on its applications in LAE. Furthermore, we present a case study using a generative diffusion model (GDM)-based algorithm to enhance the performance of aerial collaborative beamforming-enabled remote secure communications in LAE and simulation results verified the effectiveness of the proposed algorithms. Finally, promising research opportunities are identified.
@@ I Introduction
Fortunately, recent advances in artificial intelligence (AI) technologies offer promising solutions to these issues. Traditional AI (TAI) techniques primarily target discriminative tasks such as classification and regression, where the main objective is to map inputs to predefined outputs. In the context of beamforming optimization, TAI techniques typically map observed channel conditions directly to predefined optimal beamforming parameters. Although these TAI methods perform well for classification or regression tasks in known scenarios, they inherently depend on labeled datasets and fixed optimization objectives. In contrast, generative artificial intelligence (GenAI) models focus on identifying the underlying distribution of the data. By capturing the intrinsic data distributions, GenAI models establish probabilistic models that identify hidden relationships within unstructured environmental data. Consequently, GenAI methods are capable of synthesizing optimal beamforming configurations through probabilistic inference, thereby generalizing beyond historical scenarios. This shift from mapping inputs to predefined outputs towards a data-driven generative approach fundamentally improves the adaptability of dynamic networks and simultaneously reduces the reliance on labeled datasets.
@@
By applying generative models, GenAI methods can capture complex channel characteristics and generate highly accurate channel state information (CSI), thereby improving the design and adjustment strategies for beamforming. Moreover, GenAI models present potential for adapting to frequent channel variations in dynamic environments and enhancing resource utilization and interference management. Additionally, GenAI models can learn from diverse data distributions, which enables them to evolve given various communication requirements and lays the foundation for multi-scenario applications within LAE.
@@ II-A Low-Altitude Economy
LAE is an emerging economic model that uses low-altitude airspace to tackle the challenges of modern cities. By incorporating advanced communication technologies and intelligent computing into low-altitude platforms (LAPs), LAE improves the efficiency and responsiveness of industries such as logistics, urban mobility, and emergency services. The key characteristics and challenges of LAE are as follows.
II-A1 Diverse Mission Types
LAE is marked by diverse missions. For instance, LAPs are used for aerial photography, and help with property promotion. Also, in medical emergencies, LAPs transport crucial supplies. Thus, this variety calls for customized configurations and specialized operations.
II-A2 Decentralized Communication
LAE uses ad hoc communications, where aerial vehicles communicate directly with each other without a centralized controller. This enables flexible coordination and ensures reliability in environments where traditional infrastructure may be limited or unavailable [3].
II-A3 Complex Environments
Low-altitude operations are affected by urban structures, weather conditions, and obstacles. Therefore, adaptive sensing and communication strategies are essential to ensure stable and reliable operations.
@@ II-A3 Complex Environments
Given the aforementioned characteristics and challenges of LAE, effective wireless signal management and optimization in low-altitude operations are crucial. Hence, beamforming emerges as a relevant solution, which has substantial potential in tackling signal-related challenges.
@@
II-B Beamforming
Beamforming is a signal processing technique for optimizing the directionality of electromagnetic waves by controlling the phase and amplitude of each antenna element in the array antenna. Fig. 1 illustrating the key technologies involved in beamforming is provided, and the details of these technologies are as follows.
II-B1 Beam Management
Beam management focuses on acquiring and maintaining reliable beam pairs to ensure robust directional communication. In LAE, it maintains stable links for certain applications such as drone logistics and aerial monitoring, which ensures uninterrupted communication for mission-critical tasks. To achieve this, beam management includes four key steps, which work together to maintain optimal communication paths in real time. These steps are as follows.
  • Beam Sweeping: Beam sweeping involves scanning predefined beam directions to identify optimal signal paths. In LAE, rapid beam sweeping addresses the high mobility of LAPs, thus enabling quick alignment with ground nodes or other aerial devices.
  • Beam Measurement: Beam measurement evaluates signal quality metrics including strength and interference across candidate beams. In LAE, this process dynamically adapts to environmental fluctuations such as weather changes or urban reflections to sustain robust link quality.
  • Beam Determination: Based on the results of beam measurements, beam determination selects the optimal beam direction and width. This step is critical for LAE to prioritize low-latency links for time-sensitive tasks, such as collision avoidance or real-time surveillance.
  • Beam Reporting: Beam reporting involves transmitting feedback on beam status to network controllers or centralized systems. This process supports coordinated multi-LAP operations, thereby minimizing interference and optimizing spectrum utilization.
II-B2 Beam Tracking
Beam tracking dynamically adjusts the beam direction in real time to follow the movement of targets. In LAE operations, where targets such as moving vehicles or other LAPs frequently change their positions, beam tracking helps LAPs maintain stable communication links.
II-C Applications of Beamforming in Low-Altitude Economy
II-C1 Communications
Beamforming plays a critical role in LAE communications. First, beamforming reduces interference by directing signal energy more accurately toward a specified target through the enhanced directivity of the antenna array. This advantage becomes especially significant in dense urban environments and large-scale events, where multiple signal sources compete for limited spectrum resources [4]. With another aspect of beam focusing, beamforming strengthens security by making signal interception significantly more challenging for unintended parties. Moreover, beamforming enhances spectrum utilization via effective interference mitigation by tailoring beams considering both the direction and distance according to the location of each user or device, which enables more efficient spectrum reuse and allows multiple users to share the same frequency band simultaneously.
II-C2 Sensing
Beamforming can effectively address the sensing tasks in LAE applications. First, beamforming enhances the ability of LAPs to track targets by adjusting the beam direction based on the real-time position of the target, which ensures consistent focus and maintains tracking accuracy even with changing trajectories. Furthermore, beamforming strengthens the ability of LAPs to monitor the environment. Specifically, by focusing wireless signal energy, beamforming increases signal strength in target areas, which helps resolve attenuation during long-distance transmission [5].
@@ II-D1 Scalability Constraints
II-D2 Real-Time Adaptation
@@
II-D3 Multi-Objective Optimization
Beamforming balances the interference reduction, energy efficiency, and communication reliability. However, it often faces stringent environmental and resource constraints. As a result, overlapping paths and limited spectrum in LAE scenarios make emergency communications and similar high-demand situations increasingly challenging. Therefore, addressing these multi-objective challenges calls for more advanced optimization techniques.
II-D4 Antenna Technology Complexity
@@
TABLE I: The Use of Generative AI in Beamforming and Beamforming in Low-Altitude Economy.
Issues Generative AI Model Potential Benefits
GAN VAE GDM Transformer
Beam Management Simulate robust feedback channels Denoise multimodal inputs Optimize beam direction through probabilistic generation and refinement Identify relationships across modalities [6] 1) Improve the accuracy of beamforming 2) Increase the real-time responsiveness of beamforming 3) Reduce the computational overhead of beamforming
Beam Tracking Generate diverse environmental training datasets [7] Encode channel features into latent space Enable parallel tracking with multi-head attention
Beamforming in LAE Improve channel modeling and estimation [8] [9] and reduce energy Extract features to generate beam and spectrum allocation strategies Generate adaptive beam directions through learned patterns Capture long-range dependencies and model context [10] 1) Enhance channel characterization and estimation 2) Optimize energy consumption, resource allocation, and security
@@ III-A Fundamentals of Generative AI
III-A1 Generative Adversarial Network (GAN)
GANs are a class of GenAI models that consist of two competing neural networks, which are widely used in areas such as image generation and data augmentation. The core principle of GANs lies in adversarial training between the generator and the discriminator, where the generator progressively enhances the realism of the generated data while the discriminator improves its ability to distinguish real data from fake. In wireless communications, GANs contribute to mitigating noise by modeling complex data distributions, thereby leading to more stable and robust communication. For instance, in [11], a GAN-based model was designed to improve bit error rate (BER) performance in uplink Rayleigh fading channels. As a result, the decreased BER improves the accuracy of channel state information (CSI) estimation. This then allows for better adjustment of beamforming weights to focus energy on target users while suppressing interference, which increases the signal-to-interference-plus-noise ratio (SINR). Moreover, the enhanced SINR facilitates more accurate CSI estimation through cleaner feedback. During this process, beamforming achieves precise signal alignment by dynamically adjusting phase and amplitude, performing matched filtering through the channel to maximize array gain, ultimately leading to improved overall communication performance.
III-A2 Variational Autoencoder (VAE)
Apart from GANs, VAEs represent another major class of GenAI models that learn latent representations of input data and generate similar outputs by maximizing the data likelihood. Each model consists of an encoder that compresses the data into latent parameters, and a decoder that reconstructs the data from this distribution. In wireless communications, the unique encoding and decoding mechanisms of VAEs enhance information security. For example, in [12], the authors proposed a VAE-based authentication scheme that encodes channel impulse response features into a probabilistic latent space, enabling secure authentication without requiring the channel prior knowledge of attacker. Thus, VAEs can support secure beamforming by encoding channel features into latent distributions, dynamically optimizing parameters to resist interference.
III-A3 Generative Diffusion Model (GDM)
In addition to GANs and VAEs, GDMs have emerged as a powerful class of GenAI models based on probabilistic denoising processes, which generate data through forward and reverse diffusion processes. Specifically, the forward process gradually adds noise to transform the data into a Gaussian distribution, while the reverse process removes the noise to reconstruct the data and recover the target distribution. In wireless communications, GDMs are valuable for channel simulation, as they can replicate complex channel characteristics. For example, in [13], the authors applied a GDM-based model to generate channel data conditioned on user positions, producing realistic MIMO channel samples even with limited real measurements. Consequently, beamforming algorithms can leverage these simulated channel conditions to improve beam direction estimation, particularly in sparse channel measurement scenarios.
III-A4 Transformer
Transformer models are deep learning architectures centered around attention mechanisms and have gained prominence in various generative tasks. In this paradigm, input data representing temporal and spatial variations in network behavior is first converted into tokens. Unlike GANs, VAEs, and GDMs, which generate samples by sampling from a learned latent distribution, Transformer-based models generate content by modeling the conditional probability of each element in a sequence, producing outputs in either an autoregressive or parallel manner. In wireless communications, Transformer models can facilitate channel estimation tasks. For example, in [14], the authors embedded dynamic network states into structured tokens and employed a Transformer to capture multi-scale temporal and spatial dependencies. The resulting representations effectively modeled time-varying channel dynamics and user preferences, thereby enhancing the exploration performance of actor-critic networks in deep reinforcement learning. As such, Transformer-based approaches can optimize beamforming by dynamically adjusting beam parameters in response to real-time network conditions.
III-B Generative AI for Beamforming
III-B1 Generative AI Solutions in Beam Management
Conventional beam management approaches face notable limitations in dynamic and complex environments. First, conventional beam management approaches employ oversimplified beam prediction models that lack the adaptability to rapid channel variations caused by mobility and environmental obstructions. Additionally, these methods primarily utilize single-modality data while ignoring multimodal contextual information from supplementary sensing inputs like visual perception systems. Furthermore, rigid spatial coordination strategies and linear optimization frameworks fail to capture the long-range spatiotemporal correlations, resulting in suboptimal beam alignment and compromised throughput efficiency. In contrast, GenAI-enhanced methods address these limitations by leveraging multi-modal data fusion and adaptive learning.
For instance, in [6], the authors proposed a multi-modal Transformer model combined with reinforcement learning to optimize beam management. Specifically, the model processes visual, light detection and ranging (LiDAR), and radar data by using self-attention mechanisms to identify the relationships across modalities. This enables efficient prediction of optimal beam groups by reducing the computational complexity through narrowed search spaces.
Beyond holistic beam management, GenAI techniques can further enhance specific sub-processes. For beam sweeping, GDMs could synthesize diverse environmental scenarios such as varying pedestrian densities or weather conditions to optimize initial beam scans. In beam measurement, VAEs can denoise LiDAR or radar inputs to enhance signal quality and ensure reliable beam measurements. Furthermore, the feature fusion capability of multi-modal Transformer directly aids beam determination by integrating multi-modal data to prioritize high-SINR beams. Finally, for beam reporting, GANs could simulate realistic feedback channels, thus ensuring robust reporting under latency or packet loss.
III-B2 Generative AI Solutions in Beam Tracking
Conventional beam tracking algorithms require real-time calculations of the movement of target devices and channel states to adjust the beam direction. However, in complex environments with interference and blockages, rapid changes in channel conditions can make it difficult for tracking to stay aligned with the target. Additionally, conventional methods often rely on simple motion models, such as linear or Kalman filter-based predictions, which struggle to accurately predict the highly uncertain movement trajectories of user devices. In contrast, GenAI models can generate comprehensive training datasets covering various environmental scenarios and complex movement trajectories.
For instance, in [7], the authors proposed a GAN-based method to generate diverse and comprehensive training datasets. These datasets include variations such as different human postures, environmental changes, as well as multipath and blockage scenarios. Furthermore, such datasets can be used to enhance the training of deep learning models, thereby improving their ability to adapt to rapidly changing environments and ensuring more stable and accurate beam tracking.
III-C Generative AI for Beamforming in Low-Altitude Economy

### v2

@@ Abstract
The growth of low-altitude economy (LAE) has driven a rising demand for efficient and secure communication. However, conventional beamforming optimization techniques struggle in the complex LAE environments. In this context, generative artificial intelligence (GenAI) methods provide a promising solution. In this article, we first introduce the core concepts of LAE and the roles of beamforming in advanced communication technologies for LAE. We then examine their interrelation, followed by an analysis of the limitations of conventional beamforming methods. Next, we provide an overview of how GenAI methods enhance the process of beamforming, with a focus on its applications in LAE. Furthermore, we present a case study using a generative diffusion model (GDM)-based algorithm to enhance the performance of aerial collaborative beamforming-enabled remote secure communications in LAE and simulation results demonstrate that our approach achieves 23% improvement in terms of secrecy rate and 18% reduction in energy consumption compared to baseline algorithms. Finally, promising research opportunities are identified.
@@ I Introduction
Fortunately, recent advances in artificial intelligence (AI) technologies offer promising solutions to these issues. Traditional AI (TAI) techniques are built on supervised learning, directly mapping labeled channel data to optimal beamforming parameters. Moreover, TAI-based methods utilize deterministic algorithms with reduced parameter dimensionality and memory footprint. As a result, they are well suited for the lightweight deployment, which requires minimal computational resources for implementation on resource-constrained low-altitude platforms (LAPs). However, the reliance of TAI on predefined mappings and labeled datasets limits its generalization and adaptability in highly dynamic LAE environments with non-stationary channels or sparse data.
In contrast, generative AI (GenAI) models can inherently address these challenges. Specifically, GenAI models use the self-supervised learning enables diverse environment adaptation by capturing stochastic propagation phenomena and transferring knowledge across varied conditions, providing robust performance without environment-specific tuning. Furthermore, GenAI can synthesize realistic channel realizations and adapting to new situations with minimal training data. Additionally, uncertainty quantification mechanisms in generative models support optimal decision-making under limited channel state information, effectively balancing performance with data constraints in challenging LAE communication scenarios.
@@
By applying generative models, GenAI methods can capture complex channel characteristics and generate highly accurate channel state information (CSI), thereby improving the design and adjustment strategies for beamforming. Moreover, GenAI models present potential for adapting to frequent channel variations in dynamic environments and enhancing resource utilization and interference management. Crucially, GenAI models can learn from diverse data distributions, which enables them to evolve given various communication requirements and lays the foundation for multi-scenario applications within LAE.
@@ II-A Low-Altitude Economy
LAE is an emerging economic model that uses low-altitude airspace to tackle the challenges of modern cities. By incorporating advanced communication technologies and intelligent computing into LAPs, LAE improves the efficiency and responsiveness of industries such as logistics, urban mobility, and emergency services. The key characteristics and challenges of LAE are as follows.
II-A1 Heterogeneous Communication Demands
The diverse operational requirements of LAPs in LAE scenarios create heterogeneous communication demands, thus bringing significant challenges in communication resource allocation. Specifically, in disaster reconnaissance and urban surveillance scenarios, LAPs must support high data rates to enable real-time video transmission [3]. In comparison, applications such as urban air taxis rely on ultra-reliable links with extremely low latency, as both reliability and responsiveness are critical for ensuring passenger safety. Meanwhile, for agricultural protection, LAPs prioritize wide coverage over transmission rate. Consequently, these varied and often conflicting requirements necessitate advanced communication solutions capable of dynamic resource allocation.
II-A2 Dynamic Network Topologies
The mission-driven mobility of LAPs creates rapidly shifting network structures that challenge conventional communication paradigms. Specifically, large-scale agricultural spraying operations need real-time topology adaptation as LAP swarms cover vast farmlands [4]. Moreover, emergency response, such as wildfire monitoring, requires LAPs to dynamically form temporary relay chains for extended coverage in disaster zones. In addition, for urban air logistics, delivery LAP fleets execute sub-second topology switches for collision avoidance in congested corridors. Consequently, mission scenarios involving highly dynamic LAP mobility demand self-organizing networks capable of seamless topology transitions under strict latency constraints.
II-A3 Complex Communication Environments
LAP operations face severe wireless signal disruption. Specifically, urban package delivery encounters unpredictable obstructions and reflections from buildings and alleys. Moreover, in precision crop dusting over terraced fields, agricultural LAPs suffer signal degradation and physical obstructions from dense foliage. Furthermore, event monitoring over crowded areas experiences intense interference from numerous nearby LAPs, phones, and other wireless devices. As a result, these environmental and interference challenges necessitate robust communication strategies adaptable to rapidly changing signal conditions for reliable and stable system performance.
@@ II-A3 Complex Communication Environments
Given the aforementioned characteristics and challenges of LAE, effective wireless signal management and optimization in low-altitude operations are crucial. Under this background, beamforming emerges as a relevant solution, which has substantial potential in tackling signal-related challenges.
II-B Beamforming
Beamforming is a signal processing technique for optimizing the directionality of electromagnetic waves by controlling the phase and amplitude of each antenna element in an antenna array. To illustrate the key technologies involved in beamforming, Fig. 1 presents a timeline of representative methods. The details of these technologies are discussed below.
II-B1 Beam Management
Beam management focuses on acquiring and maintaining reliable beam pairs to ensure robust directional communication.
  • Beam Sweeping: Beam sweeping involves scanning predefined beam directions to identify optimal signal paths.
  • Beam Measurement: Beam measurement evaluates signal quality metrics including strength and interference across candidate beams.
  • Beam Determination: Based on the results of beam measurements, the beam determination algorithm selects the optimal configuration.
  • Beam Reporting: Beam reporting involves transmitting feedback on beam status to network controllers or centralized systems.
II-B2 Beam Tracking
Beam tracking continuously updates the beam direction in response to target movement, which ensures that the beam remains accurately aligned throughout the transmission process.
II-C Beamforming Solutions in Low-Altitude Economy
II-C1 Beamforming for Heterogeneous Communication Demands
Beamforming offers an effective means of addressing heterogeneous communication demands in LAE scenarios by directing signal energy toward specific spatial directions. For instance, for high data rate applications, beamforming enhances the throughput of data-intensive transmissions by focusing strong beams on video-streaming LAPs. Conversely, in ultra-reliable communication scenarios, beamforming helps establish robust links for safety-critical platforms by precisely steering energy and suppressing potential interference. Furthermore, for wide-area coverage, beamforming extends communication range by targeting distant agricultural LAPs. Therefore, this spatial flexibility enables adaptive resource allocation tailored to the specific communication needs of each LAP, thereby facilitating the efficient coordination of diverse demands across the network.
II-C2 Beamforming for Dynamic Network Topologies
Beamforming effectively addresses the challenge of constantly changing network connections in LAE by enabling rapid redirection of communication signals. For example, in scenarios such as multi-LAP emergency response operations, beamforming allows communication nodes to quickly shift their focused signal beams as LAPs relocate or reconfigure formations. Similarly, for large-scale LAP-based delivery systems, beamforming enables central communication hubs to dynamically track and maintain robust signal links with other LAPs during critical maneuvers such as takeoff, landing, and obstacle navigation. As a result, by continuously adjusting the beam direction towards moving targets, beamforming significantly improves connection stability and minimizes disruptions in these highly dynamic environments.
II-C3 Beamforming for Complex Communication Environments
Beamforming provides a powerful solution for overcoming signal degradation in challenging low-altitude settings by concentrating signal energy along desired paths. Specifically, for LAPs operating in urban areas, beamforming helps overcome signal blockage caused by buildings by focusing the transmission towards the intended receiver, thereby finding stronger paths through or around obstacles, which improves link reliability. Moreover, in agricultural applications, beamforming effectively counters the signal weakening and physical obstruction caused by dense crops by directing a strong and focused beam, thus ensuring stable control and data transfer over longer distances. Furthermore, in crowded event areas with many LAPs and devices, beamforming minimizes interference by precisely directing signals towards intended receivers while reducing energy sent towards potential sources of interference. Consequently, this focused approach is essential for maintaining clear communication in congested and cluttered environments.
II-D Applications of Beamforming in Low-Altitude Economy
@@
II-D1 Communications
In the rapidly evolving landscape of LAE applications, beamforming technology can address the unique communication challenges encountered by aerial networks. First, LAE networks frequently contend with severe interference in dense urban environments where multiple signal sources compete for limited spectrum resources [5]. To address the challenge, beamforming effectively reduces interference by directing signal energy more accurately toward a specified target through the enhanced directivity of the antenna array. Therefore, such directional energy transmission reduces signal spread and minimizes interference with other communication links. For instance, in urban air mobility with high-density LAP traffic, beamforming ensures that energy is precisely concentrated on the intended communication nodes, thereby preventing signal bleed-over into adjacent channels or interfering with other aerial vehicles.
In addition to interference, spectrum scarcity poses another critical challenge in LAE systems. To overcome this challenge, beamforming enhances spectrum utilization by enabling efficient spectrum reuse. Specifically, by tailoring beam directions based on the location of each user or device, multiple users can simultaneously operate on the same frequency band without introducing significant co-channel interference. For example, in live event broadcasting, media LAPs can generate directional beams, which enables multiple LAPs to simultaneously transmit on a shared frequency channel to different ground receivers across a venue, thereby effectively managing co-channel interference.
II-D2 Sensing
Beamforming plays a pivotal role in overcoming key sensing challenges in LAE applications. First, one significant challenge in LAE sensing is maintaining accurate tracking of moving targets, especially those with highly dynamic and unpredictable motion patterns. To address this issue, beamforming dynamically adjusts the beam direction according to the real-time positions of the targets. Specifically, by continuously aligning signal focus with the instantaneous location of each target, beamforming supports consistent tracking precision under mobility. For instance, in LAE logistics scenarios involving LAP swarms, precise and uninterrupted target localization is essential. In such scenarios, adaptive beam control enables each LAP to maintain accurate localization even as target positions fluctuate, ultimately supporting mission reliability.
In addition to target tracking, another critical challenge is ensuring robust environmental sensing, particularly when wireless signals are prone to significant attenuation over long distances or when attempting to penetrate obstructions like debris or foliage. Moreover, such attenuation can severely degrade the quality of sensing data. To address this issue, beamforming mitigates attenuation effects by concentrating wireless signal energy toward areas of interest, thereby improving signal strength and increasing the fidelity of received data  [6]. For example, in LAE disaster response missions, concentrated beam scanning enables rapid detection of survivors hidden beneath rubble or obstructive terrain, thus enabling responders to efficiently locate individuals even in challenging environments where weak signals would render detection impossible.
@@ II-E1 Scalability Constraints
II-E2 Real-Time Adaptation
@@
II-E3 Multi-Objective Optimization
Beamforming balances the interference reduction, energy efficiency, and communication reliability. However, beamforming often faces stringent environmental and resource constraints. Consequently, overlapping paths and limited spectrum in LAE scenarios make emergency communications and similar high-demand situations increasingly challenging. Therefore, addressing these multi-objective challenges calls for more advanced optimization techniques.
II-E4 Antenna Technology Complexity
@@ III-A Fundamentals of Generative AI
III-A1 Variational Autoencoders (VAEs)
VAEs establish a probabilistic generative framework by explicitly optimizing a lower bound on data likelihood to learn latent representations that capture underlying data structures. Their distinctive advantage lies in naturally quantifying uncertainty through explicit variance modeling, while simultaneously ensuring consistent training convergence, despite typically generating reconstructions with less sharp details. As a result, the inherent uncertainty quantification makes VAEs particularly valuable for channel estimation in rapidly changing LAE conditions, where accurate beam prediction requires modeling the distribution of possible channel states rather than providing single point estimates. In light of these properties, the authors in [7] employed a VAE-based model to learn the complex probability distribution of CSI, thus parameterizing estimators that approach minimum mean square error performance even without perfect CSI during training phases. Moreover, by encoding channel characteristics into a latent space with explicit variance parameters, VAEs provide not only the channel estimates themselves but also confidence measures for those estimates, proving crucial when making communication decisions in variable channel conditions. In addition, the robustness of VAEs stems from this probabilistic generative structure. Specifically, by modeling a distribution over latent variables, VAEs inherently account for uncertainty and variability, thus promoting strong generalization and robust inference from partial, noisy, or degraded input, which is beneficial when sensor data is missing or corrupted.
III-A2 Generative Diffusion Models (GDMs)
GDMs learn complex data distributions through a distinctive iterative denoising process. Specifically, GDMs simulate a gradual forward diffusion process that incrementally adds noise to data. Subsequently, the models learn to precisely reverse this process by progressively removing the noise step by step, thereby reconstructing the original data. Therefore, the step-wise denoising mechanism enables GDMs to capture the full transformation from pure noise to structured data, along with underlying statistical structures, thus leading to the generation of exceptionally high-fidelity and diverse samples. Moreover, this remarkable ability to generate accurate and varied data makes GDMs particularly valuable for modeling complex LAE channels, especially when detailed real-world channel measurements are limited or challenging to obtain. In light of these properties, in [8], the authors applied a GDM-based model to generate realistic MIMO channel samples conditioned on user location, even with limited real-world measurements. In addition, the robustness of GDMs is rooted in this iterative denoising procedure, thereby allowing them to correct arbitrary levels of corruption and achieve resilient reconstruction under strong jamming, complex noise, or signal dropout conditions, all common in low-altitude environments [9].
TABLE I: Generative AI Models for Beamforming Applications in Low-Altitude Economy
Model Channel Modeling & Estimation Resource Allocation Communication Security Energy Consumption Real-time Performance
VAE Quantify uncertainty in channel estimation by modeling probability distributions [7] Extract features to generate beam and spectrum allocation strategies - - Text: Moderate Image: Moderate Video: Moderate
GDM Generate realistic MIMO channel samples via iterative denoising process [8] - Resist eavesdropping interference and mitigate location errors in beamforming [10] Iteratively refine beam patterns for energy-efficient power distribution Text: Fair Image: Fair Video: Fair
GAN Improve channel modeling and estimation [11], [12] - Suppress complex interference patterns through adversarial denoising [13] Simulate low-power beam patterns Text: Excellent Image: Excellent Video: Good
Transformer Identify relationships across modalities and enable parallel tracking for estimation Capture long-range dependencies and context for resource allocation Enhance secrecy via beamforming and spatiotemporal modeling [10] Optimize energy consumption by capturing environmental dependencies Text: Good Image: Good Video: Moderate
III-A3 Generative Adversarial Networks (GANs)
GANs learn data distributions through an adversarial training framework, where a generator and a discriminator are trained in competition to implicitly capture underlying data structures. Crucially, the key strength of GANs lies in their ability to learn complex data patterns without needing explicit mathematical modeling of the data distribution. Therefore, the capability is particularly advantageous for interference suppression in LAE, which is characterized by non-stationary and complex interference patterns. Moreover, by training on clean signal data, GANs can learn what constitutes a clean signal and then use this knowledge to effectively remove interference from corrupted signals. As a result, GAN-based architectures enables robust interference mitigation even when the interference characteristics are unknown or constantly changing. For example, in [13], the authors proposed a GAN-based model for wireless signal denoising where the generator reconstructed clean signals from noisy inputs and the discriminator ensured that the outputs were indistinguishable from real clean signals, significantly improving clarity under complex noise and interference. In addition, the robustness of GANs arises from these adversarial training dynamics. Specifically, continuous competition forces the generator to approximate data distributions, including rare or extreme variants, which promotes strong resilience to distributional shifts and adversarial perturbations, frequently occurring in dynamic wireless environments.
III-A4 Transformers
Transformers leverage global context through self-attention mechanisms, which generate content by modeling the conditional probability of each element in the sequence. A key advantage of self-attention lies in its ability to directly capture long-range dependencies and relationships within sequential data, rather than relying on fixed-size windows or recurrent connections. Consequently, the inherent capability for modeling comprehensive spatiotemporal relationships makes Transformers particularly valuable for tasks requiring strong temporal and spatial awareness. For example, in [14], the authors embedded dynamic network states into structured tokens and applied a Transformer-based model to model multi-scale spatiotemporal dependencies. The resulting representations effectively captured time-varying channel dynamics and user preferences, thereby enhancing the exploration capability of the actor-critic network in deep reinforcement learning (DRL). In addition, the robustness of Transformers derives from the self-attention mechanism and their capacity to process non-sequential, variable-length, and partially corrupted input. Specifically, self-attention enables the model to dynamically reweight input tokens based on contextual relevance, thus allowing suppression of noisy features and inference of missing information by referencing global context. Consequently, the robustness of Transformers makes them particularly well suited for handling irregular input streams, fluctuating channel quality, and asynchronous data transmissions in LAE scenarios.
III-B Generative AI for Beamforming in Low-Altitude Economy